OpenStudy (anonymous):

please helllp! how to find a quadratic function given the maximum value of 6 when x=-2 and one zero is 1?

OpenStudy (anonymous):

you can start by writing it in vertex form as \[f(x)=a(x+2)^2+6\]

OpenStudy (anonymous):

then solve for \(a\) by setting \(f(1)=0\)

OpenStudy (anonymous):

Why do I set it to zero?

OpenStudy (anonymous):

Or how do I know to set it to zero?

jimthompson5910 (jim_thompson5910):

`one zero is 1` means that f(1) = 0 `one zero is 1` is another way of saying `x = 1 is a root of f(x)` or `x = 1 is an x intercept of f(x)`

OpenStudy (anonymous):

\[f(1)=0\\ f(1)=a(1+2)^2+6=0\]

OpenStudy (anonymous):

i.e. \[9a+6=0\] solve that one

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

whoaa wait hold the phone is there a typo in this question?

OpenStudy (anonymous):

" find a quadratic function given the minimum value of 6 when x=-2 and one zero is 1?"

OpenStudy (anonymous):

maximum...

OpenStudy (anonymous):

or does it say maybe "MAXIMUM"?

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

and yes \[a=-\frac{2}{3}\]